Positive charge of 2 micro coulomb and 8micro coulomb are placed 15 centimeter apart at what distance from smaller charge will the electric field due to them be zero

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
E = k Q/r²
Hence
E1 = k*2/x²
And 
E2 = k*8/(15-x)²
2k/x² = 8k/(15-x)²
15-x = 2x
3x = 15
Hence
X = 5 cm from smaller charge
 
Regards
Arun (askIITians forum expert)